OpenStudy (anonymous):

help pls Multiply and Simplify (z3-64) / (z2-5z+4) * (2z2+9z-11) / (5z2+20z+80)

OpenStudy (dumbcow):

try to factor each part, then you can cancel

OpenStudy (dumbcow):

\[z^{3}-64 = z^{3} -4^{3} = (z-4)(z^{2}+4z+16)\] \[z^{2}-5z+4 = (z-4)(z-1)\] \[2z^{2}+9z-11 = (2z+11)(z-1)\] \[5z^{2}+20z+80 = 5(z^{2}+4z+16)\]

OpenStudy (anonymous):

so then do I cancel out the ones that are the same

OpenStudy (dumbcow):

yes since you are multiplying, if same term is both on top and bottom they cancel

OpenStudy (anonymous):

leaving (2z+11) / 5

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

woo hoo thank you

OpenStudy (dumbcow):

:)